write the standard form of the equation of the parabola with its vertex (0,0)and focus at (0,5).
Answer by DSMLMD(16)

Vertex (0,0) and Focus at (0,5)


Vertex:
(0,0)
(a,b)


Focus Point:
(0,5)
(0,p)


we get p = 5


So because the point of focus located in y-coordinate, so the parabola equation is:


(x - a)^2 = 4p(y - b)
(x - 0)^2 = 4(5)(y - 0)
x^2 = 20y


The parabola equation is x^2 = 20y or x^2 - 20y = 0